### v1

@@ 2.7 YOLO-CIANNA complete loss function
Depending on the application, it might be useful to balance the relative importance of the predicted quantities. For example, high precision on the predicted size might not be necessary when detecting small objects, while center position becomes the main factor of prediction quality. Following the classical YOLO approach, we have loss scaling factors for the box position $\lambda_{pos}$, the box size $\lambda_{size}$, the probability $\lambda_{prob}$, the objectness $\lambda_{obj}$, the classes $\lambda_{class}$, and the extra parameters $\lambda_{param}$.
While loss scaling balances the general final importance of each predicted quantity, it does not allow the network to guide its expressivity regarding its current prediction dynamically. For example, adjusting the predicted class of a detection unit that does not yet properly position or detect the corresponding object is irrelevant, and it might even reinforce wrong features or add noise to the training. To add this capability, we enriched the custom loss of our YOLO-CIANNA method with prediction quality limits over some predicted properties. While the position and the size are always updated for a match, we added quality limits for the objectness $L^{{\rm fIoU}}_{O\_good}$, the probability $L^{{\rm fIoU}}_{P\_good}$, the classification $L^{{\rm fIoU}}_{C\_good}$, and the extra-parameters $L^{{\rm fIoU}}_{p\_good}$. Each loss sub-part is set to zero if the current ${\rm fIoU}$ between the target and predicted boxes is below its specific quality threshold.
This quality limit principle, combined with other association refinements, results in what we call a ”cascading loss” that varies during training to guide the network expressivity toward the important aspects, not adjusting currently irrelevant properties. A complete description of this process and the effect it has on training performances and loss monitoring is given in Appendix A.7. With all the previously introduced elements, we can define our complete YOLO-CIANNA loss function for one image as
$\begin{split}\mathcal{L}=&\sum_{i=0}^{N_{g}}\sum_{j=0}^{N_{b}}\mathds{1}_{ij}^{match}\begin{aligned} \Biggl{(}&\lambda_{pos}&&\left[(o^{x}_{ij}-\hat{o}^{x}_{ij})^{2}+(o^{y}_{ij}-\hat{o}^{y}_{ij})^{2}\right]\\ +&\lambda_{size}&&\left[(o^{w}_{ij}-\hat{o}^{h}_{ij})^{2}+(o^{h}_{ij}-\hat{o}^{h}_{ij})^{2}\right]\\ +&\lambda_{class}&&\mathds{1}_{ij}^{C\_good}\sum_{k}^{N_{C}}\left(-\hat{C}_{ij}^{k}\log(C_{ij}^{k})\right)\\ +&\lambda_{param}&&\mathds{1}_{ij}^{p\_good}\sum_{k}^{N_{p}}\gamma^{k}\left(p^{k}_{ij}-\hat{p}^{k}_{ij}\right)^{2}\\ +&\lambda_{prob}&&\mathds{1}_{ij}^{P\_good}\left(P_{ij}-1\right)^{2}\\ +&\lambda_{obj}&&\mathds{1}_{ij}^{O\_good}\left(O_{ij}-{\rm fIoU}_{ij}\right)^{2}\Biggl{)}\end{aligned}\\ +&\sum_{i=0}^{N_{g}}\sum_{j=0}^{N_{b}}\mathds{1}_{ij}^{void}\lambda_{void}^{j}\begin{aligned} &\Biggl{(}\lambda_{prob}\left(P_{ij}-0\right)^{2}+\lambda_{obj}\left(O_{ij}-0\right)^{2}\Biggl{)}.\end{aligned}\\ \end{split}$ (14)
In this equation, in addition to the already defined parameters, the first sum runs over all the elements of the output grid for a single image $N_{g}$, and the second sum runs over all the detection units in a grid cell $N_{b}$. All the values with a hat represent the targets for the corresponding predicted values, and the $\mathds{1}_{ij}^{X\_good}$ are masks that identify the predicted boxes that respect the prediction quality limit of each sub-parts. This loss is written for a classification based on a soft-max activation with a cross-entropy error, but a sigmoid activation with a sum-of-square error is also available in our implementation.
@@
Figure 5: Illustration of the NMS process in a given image. The dashed boxes are the target, and the solid boxes are the predictions. The line widths of the boxes are scaled on their respective objectness score. The colors indicate the state of the box in the NMS process at different steps. Frame (a) represents the targets and the raw network predictions that remain after the objectness filtering. Frames (b) and (c) represent two successive steps of the NMS process with a different best current box. Frame (d) represents the remaining boxes after completion of the NMS.
@@ 2.8 Prediction filtering and non-maximum suppression
We expect a properly trained detector to order its predictions by quality based on the predicted objectness score for each detection unit. The raw detector output is always a static list of boxes of size $g_{w},g_{h},N_{b}{\times}(6+N_{c}+N_{p})\right>$ regardless of the input content. Consequently, the predicted boxes must be filtered based on their objectness score to remove those that are unlikely to represent an object. By design, the number of actually detectable objects in the image should be low compared to the total number of detection units in the grid. Therefore, most of the predicted boxes belong to the background type with a low objectness score.
@@
While the continuous objectness score is the best direct representation of the detector’s inner workings, it is incompatible with some final metric that needs a list of considered ”good” detections. Visualizing the predicted boxes also requires filtering to preserve only the plausible detection. In such cases, an objectness threshold can be used to remove low-confidence detections. This is usually done on a validation or test dataset not used for training but for which the targets are known. The threshold can be optimized to maximize a detection metric on this test dataset.
@@
The main difficulty is that all detection units have fitted their objectness independently. Due to the fully convolutional structure of the network, a given detection unit represents the same objectness fitting over the full grid, meaning that the same threshold can be used. On the contrary, the predicted objectness between two independent detection units is not comparable as it depends on the type and frequency of targets associated with each of them at training time. The classical solution is to fit an individual objectness threshold for each of them. However, it considers that predictions from different detection units are independent, which is not true for most applications. Still, fitting a per-detection-unit objectness threshold removes the vast majority of false positives. To achieve the best results, the objectness regimes must be homogenized between the different detection units from the start by adjusting the individual $\lambda_{void}^{b}$ factors (Sect. 2.5). While this is mainly an empirical and iterative process, the main principle is to balance the ratio between detection and background cases based on how the objects from the training sample are expected to distribute over the detection units.
@@
With most false positives removed, there can still be multiple high-objectness predictions that represent the same object. To preserve only the best-detected box for each object, we use a classical post-processing step called non-maximum suppression (NMS, Felzenszwalb et al., 2010; Girshick et al., 2013). It consists of an iterative search for the box with the highest objectness score in the image to remove all the overlapping predicted boxes. To consider that there is an overlap, the two boxes must verify ${\rm fIoU}>L^{{\rm fIoU}}_{NMS}$ with the ${\rm fIoU}$ being computed between the two predicted boxes. All boxes with a lower objectness score than the highest-scoring box are removed, and the best box is added to a static prediction list. This process is repeated until no boxes are left in the raw-prediction list. It is illustrated in Fig. 5.
@@
The NMS is done regardless of what detection unit generated the predicted boxes, demonstrating that they are not independent as they can remove each other based on their respective objectness. This is one of the main reasons to force all detection units to have similar objectness distributions. The detection quality can only be evaluated after the NMS, so searching for the best $\lambda_{void}^{b}$ factors is dependent on the $L^{{\rm fIoU}}_{NMS}$ and respectively.
@@ 3 Dataset description and network training
In this section, we present the main aspects of the SDC1 data along with the expected products and the associated metrics. A complete description of the SDC1 challenge and its data products can be found in Bonaldi et al. (2021), while the underlying T-RECS simulation is detailed in (Bonaldi et al., 2019). We also present the pre-processing of the data to construct our training sample. From this, we describe our best-performing network architecture and specify the corresponding setup and hyper-parameters for our YOLO-CIANNA detector.
3.1 Sub-challenge definition
@@
The SDC1 is a source detection and characterization task in simulated SKA-like data products (Sect. 1) that comprises nine 4GB images (three frequencies, with three integration times each) of the same field. The SDC1 is only modestly challenging regarding data volume, especially compared to the SDC2 with its 1TB data cube. Still, it represents significant challenges for detection methods in many other aspects. All the images have the same size of 32768 square pixels. As the frequency increases, the angular resolution improves while the field of view reduces. Therefore, images at different frequencies only partially overlap, meaning that the problem to solve varies with the position in the field. In addition, the number of detectable sources varies significantly with the integration time and frequency (Table 2 in Bonaldi et al., 2021). We also note that all the images are considered noise-limited, even at the highest 1000h integration time. As a benchmark for our YOLO-CIANNA, we chose to use the 560 MHz - 1000h image due to its larger field of view, higher total number of sources, and larger source density per square pixel. This image is an appropriate benchmark for our YOLO-CIANNA method, representing a difficult astronomical detection context with a high dynamic and a high density of small sources with occasional blending. While our method could technically work for the other SDC1 images, it would not be more informative regarding its capabilities. The consequences and limits of this choice are further discussed in Sect. 5.2.1.
@@
In the original challenge, participating teams were provided a True catalog for a small fraction of the image. This was supposed to facilitate method development and tuning but also allow training of supervised ML approaches. At the time, the full underlying True catalog was unavailable, and the teams had to submit their result to a remote scoring service. After the challenge, the organizers released the full True catalog, the scoring code in open source [^2] https://gitlab.com/ska-telescope/sdc/ska-sdc (Clarke & Collinson, 2021), and the submitted source catalogs from the participating teams. This allows anyone to either reproduce the challenge conditions or to use the full simulated field for other purposes. In this study, we use only the training catalog to constrain our detector to benchmark our method comparatively to other approaches. We then use the full True catalog only to present an in-depth analysis of our detector performances. Since the scorer allows individual image scoring (for all frequencies and exposure), we can produce a detection catalog for the 560 MHz - 1000h image that is directly comparable to other team submissions during and after the challenge.
@@ 3.2 Image and source catalog description
The 560 MHz - 1000h image has a field of view of $5.5$ square degrees, which contains the primary beam out to the first null. It has a full size of 32768 square pixels, a pixel size of $0.6$ arcsec, and an imaging resolution in the Gaussian approximation of $1.5$ arcsec full width at half-maximum. The noise level at this exposure and frequency is estimated at $2.55{\times}10^{-7}$ $\mbox{Jy}.\mbox{beam}^{-1}$. An example sub-field of 512 square pixels from this image is presented in Fig. 1. The simulated image contains no systematic instrumental effects like calibration, pointing, or deconvolution errors, making it unrealistically clean (Sect. 5.3). The image is not primary-beam corrected, but the corresponding primary beam image is available as ancillary data. With this setup, the instrument sensitivity decreases on the edges, but it preserves a uniform noise level over the full field. While we could convert the image to a beam-corrected one, it is better for the detector to work on the constant noise image and to detect sources from their apparent brightness (Sect. 3.5). The apparent flux of a source can be obtained by multiplying the flux from the True catalog by the interpolated primary beam value at the source’s central position.
The SDC1 challenge task is to detect and characterize the sources. The expected parameters for each source are the coordinates ($RA$, $Dec$), the integrated source flux $f$, the core fraction $c_{f}$ if different from zero, the major and minor axis ($Bmaj,Bmin$), the major axis position angle $PA$, and a classification $C$ (one of AGN-steep, AGN-flat, or star-forming galaxy). The provided True catalog provides all these properties for each source, allowing the training of supervised methods. From Sect. 5.3 in Bonaldi et al. (2021), the classification is considered difficult as it is only feasible on the fraction of the field where all frequencies are available. For this reason, we focused on getting the best performances on a detection and characterization problem only.
@@ 3.3 SDC1 scoring metric
The first element of the scorer is a match criterion. Due to source density, relying only on the central positions of the sources to establish proper matches is not enough. The SDC1 scorer uses a combination of the position, the size, and the predicted flux accuracies to represent a global matching score defined as
$E_{tot}$ $=\sqrt{E^{2}_{pos}+E^{2}_{size}+E^{2}_{flux}}\text{, with}$ (15)
$E_{pos}$ $\propto\sqrt{(x-\hat{x})^{2}+(y-\hat{y})^{2}}/\hat{S}^{\prime}\text{,}$ (16)
$E_{size}$ $\propto\left|S-\hat{S}\right|/\hat{S}^{\prime}\text{, and}$ (17)
$E_{flux}$ $\propto\left|f-\hat{f}\right|/\hat{f},$ (18)
where the true values from the catalog are indicated with a hat, ($x,y$) is the central coordinates in pixels corresponding to ($RA,Dec$), $S$ is the average value of $Bmaj$ and $Bmin$, $S^{\prime}$ is the largest axis convolved with the synthesized-beam, and $f$ is the source flux. We note that the difference between $S$ and $S^{\prime}$ is present in the scorer code but not specified in Bonaldi et al. (2021) or the challenge documentation. To prevent false detections due to the high source density, the scorer defines a strict match threshold value for $E_{tot}$. Each sub-part of the error is normalized to a value considered representative of a $1\sigma$ error. In the scorer, the normalization coefficients are set to $0.93$ for the position, $0.36$ for the flux, and $4.38$ for the size, all using the catalog units defining a global $1\sigma$ error. These individual limits were obtained by fitting the distribution of the corresponding values on the combination of all the submitted catalogs at the challenge time. A match is then defined if $E_{tot}<5\sigma$, which was optimized to reduce the average random association chance of all submitted catalogs below 10 percent. We note that the scorer distinguishes false detection into two categories: either ”False” if there is no target source closer than 1.5 times the predicted size convolved with the beam, or ”bad” if it passes this distance limit but has a too high $E_{tot}$.
From a list of identified matches, a global score that captures both the detection and the characterization has to be defined. The SDC1 scorer attributes a score of up to one point for each match depending on its characterization quality and penalizes every false detection with a strict minus one point. The characterization evaluation is decomposed into seven individual sub-scores that all respect the following scoring rule (based on the scorer code) for a single source
@@
where $j$ represents the sub-score part and $i$ a single source from the match list, $T^{j}$ is a threshold for this sub-score part, and finally, $E_{i}^{j}$ and $s_{i}^{j}$ correspond to the error term and the final sub-score part for this source. We list all the sub-score parts and their corresponding $E^{j}$ function and $T^{j}$ values in Table 1, and we illustrate the typical score as a function of the error regarding a given threshold value in Fig. 6.
We note that all of the sub-part error functions are based on a relative error, which has an asymmetric behavior. When overestimating the value, the error can rise infinitely, while underestimating a strictly positive prediction will never lower the relative error below -1. The issue is that these errors are associated with symmetric score response functions. As a consequence, the score is likely to be higher when underestimating the predictions, which is unlikely to naturally happen on quantities that have minimum practical values due to instrumental limits like the $flux$, $Bmaj$, and $Bmin$. While the effect is minor for the sub-part scoring, we note that for the matching criteria, it will result in excluding properly detected faint sources for which the noise caps the minimum predictable flux. This aspect is discussed in more detail in Sects. 4.5 and 5.1.
The final score for a given source and the average sub-part score for all sources are
$s_{i}=\frac{1}{7}\sum_{j}^{7}s_{i}^{j}\quad\text{ and }\quad\bar{s}^{j}=\sum_{i}^{N_{match}}s_{i}^{j}.$ (20)
@@
$\bar{s}=\frac{1}{N_{match}}\sum_{i}^{N_{match}}{s_{i}},$ (21)
@@
Table 1: Error functions and threshold values for all sub-scores.
Sub-part Error function $E^{j}$ Threshold $T^{j}$
@@
Major axis $\left|Bmaj-\hat{Bmaj}\right|/\hat{Bmaj}$ 0.3
Minor axis $\left|Bmin-\hat{Bmin}\right|/\hat{Bmin}$ 0.3
Position angle $\left|PA-\hat{PA}\right|$ 10.0

### v2

@@ 2.7 YOLO-CIANNA complete loss function
Figure 4: Illustration of the NMS process. Dashed boxes are the targets, and solid boxes are the predictions. The line width of each box is scaled according to its objectness score. The colors indicate the state of the box in the NMS process at different steps. Frame (a) shows the targets and the remaining detector predictions after objectness filtering. Frames (b) and (c) represent two successive steps of the NMS process with a different best current box. Frame (d) shows the selected boxes after the NMS.
Depending on the application, it might be necessary to balance the relative importance of the predicted quantities. For example, when detecting small objects, the center coordinates become the main estimate of prediction quality, while a high precision of the predicted size becomes mostly irrelevant. Therefore, we use loss scaling factors for the box position $\lambda_{\textrm{pos}}$, the box size $\lambda_{\textrm{size}}$, the probability $\lambda_{\textrm{prob}}$, the objectness $\lambda_{\textrm{obj}}$, the classes $\lambda_{\textrm{class}}$, and the extra parameters $\lambda_{\textrm{param}}$.
While loss scaling balances the general importance of each predicted quantity, it does not allow the network to guide its expressivity regarding its current prediction quality dynamically. For example, adjusting the predicted class of a detection unit that does not yet properly position or detect the corresponding object is irrelevant and can result in reinforced wrong features or noisy training. Therefore, we chose to add prediction quality limits over some predicted properties directly into the loss function. While the position and the size are always updated for a match, we added quality conditions for the objectness $L^{{\rm fIoU}}_{O}$, the probability $L^{{\rm fIoU}}_{P}$, the classification $L^{{\rm fIoU}}_{C}$, and the extra-parameters $L^{{\rm fIoU}}_{p}$. Each loss subpart is set to zero if the current ${\rm fIoU}$ between the target and predicted boxes is below its specific quality threshold.
This quality limit principle, combined with other association refinements, results in what we call a “cascading loss” that varies during training to guide the network expressivity toward the important aspects, not adjusting currently irrelevant properties. A complete description of this process and the effect it has on training performances and loss monitoring is given in Appendix A.7.
Combining all the previously introduced loss subparts, conditional masks, and scalings, we can define our complete YOLO-CIANNA loss function for one image as
$\begin{split}\mathcal{L}=&\sum_{i=0}^{N_{g}}\sum_{j=0}^{N_{b}}\mathds{1}_{ij}^{\textrm{match}}\begin{aligned} \Biggl{(}&\lambda_{\textrm{pos}}&&\left[(o^{x}_{ij}-\hat{o}^{x}_{ij})^{2}+(o^{y}_{ij}-\hat{o}^{y}_{ij})^{2}\right]\\ +&\lambda_{\textrm{size}}&&\left[(o^{w}_{ij}-\hat{o}^{w}_{ij})^{2}+(o^{h}_{ij}-\hat{o}^{h}_{ij})^{2}\right]\\ +&\lambda_{\textrm{class}}&&\mathds{1}_{ij}^{C}\sum_{k}^{N_{C}}\left(-\hat{C}_{ij}^{k}\log(C_{ij}^{k})\right)\\ +&\lambda_{\textrm{param}}&&\mathds{1}_{ij}^{p}\sum_{k}^{N_{p}}\gamma^{k}\left(p^{k}_{ij}-\hat{p}^{k}_{ij}\right)^{2}\\ +&\lambda_{\textrm{prob}}&&\mathds{1}_{ij}^{P}\left(P_{ij}-1\right)^{2}\\ +&\lambda_{\textrm{obj}}&&\mathds{1}_{ij}^{O}\left(O_{ij}-{\rm fIoU}_{ij}\right)^{2}\Biggl{)}\end{aligned}\\ +&\sum_{i=0}^{N_{g}}\sum_{j=0}^{N_{b}}\mathds{1}_{ij}^{\textrm{void}}\lambda_{\textrm{void}}^{j}\begin{aligned} &\Biggl{(}\lambda_{\textrm{prob}}\left(P_{ij}-0\right)^{2}+\lambda_{\textrm{obj}}\left(O_{ij}-0\right)^{2}\Biggl{)}.\end{aligned}\\ \end{split}$ (14)
In this equation, in addition to already defined parameters, the first sum runs over all the elements of the output grid for a single image $N_{g}$, and the second sum runs over all the detection units in a grid cell $N_{b}$. All the values with a hat represent the targets for the corresponding predicted values, and the $\mathds{1}_{ij}^{X}$ are masks of predictions that pass the quality limit of each subpart. This loss is written for a classification based on a soft-max activation and a cross-entropy error classification, but it can be modified for a sigmoid activation with a sum-of-square error.
Figure 5: Summary of the target-prediction association algorithm of YOLO-CIANNA. All steps are performed independently by each grid cell. All the elements are executed in order from top to bottom. The A, B, and C blocks represent the general case for the association. All the refinement steps of the association process corresponding to the B.2 block are described in Sect. A.6.
@@ 2.8 Prediction filtering and non-maximum suppression
We expect a properly trained detector to order its predictions by quality based on the objectness score for each detection unit. The raw detector output is always a static list of boxes of size $g_{w},g_{h},N_{b}{\times}(6+N_{c}+N_{p})\right>$ regardless of the input content. Consequently, the predicted boxes must be filtered based on their objectness score to remove those unlikely to represent an object. By design, the number of detectable objects in the image should be low compared to the total number of detection units. Therefore, most predicted boxes belong to the background type with a low objectness score.
@@
While the continuous objectness score for all the predictions is the best direct representation of the sensitivity of the detector, it is incompatible with some final metric that needs a list of considered “good” detections. Visualizing the predicted boxes also requires filtering to preserve only the plausible detection. In such cases, an objectness threshold can be used to remove low-confidence detections. The threshold is usually optimized to maximize a metric score on a validation or test dataset not used for training but for which the targets are known.
@@
Due to the fully convolutional structure of the network, objectness scores from the same detection unit can be compared over the full grid, meaning that the same threshold can be used. On the contrary, the predicted objectness between two independent detection units is not comparable as it depends on the type and frequency of targets associated with each of them during training. A possible solution is to fit an individual objectness threshold for each detection unit. However, it relies on the assumption that predictions from different detection units are independent, which is not true for most applications. In practice, it is still a good solution to remove the vast majority of false positives. To achieve the best results, the objectness regimes must be homogenized between the different detection units from the start by adjusting the individual $\lambda_{\textrm{void}}^{b}$ factors (Sect. 2.5). This is done by balancing the ratio between detection and background cases based on how the objects from the training sample are expected to distribute over the detection units.
@@
With most false positives removed, there can still be multiple high-objectness predictions that represent the same underlying object in the image. To preserve only the best-detected box for each object, we use a classical post-processing step called non-maximum suppression (NMS, Felzenszwalb et al., 2010; Girshick et al., 2013). It consists of an iterative search for the box with the highest objectness score in the image that is then used to remove all the overlapping predicted boxes. To consider that there is an overlap, the two boxes must verify ${\rm fIoU}>L^{{\rm fIoU}}_{\rm NMS}$ with the ${\rm fIoU}$ being computed between the two predicted boxes. The best box is then stored in a static list of selected detection. This process is repeated until no boxes are left in the raw-prediction list. It is illustrated in Fig. 4.
@@
The NMS is done regardless of what detection unit generated the predicted boxes, demonstrating that they are not independent as they can remove each other based on their respective objectness. This is one of the main reasons we force all detection units to have similar objectness distributions. The detection quality can only be evaluated after the NMS, so searching for the best $\lambda_{\textrm{void}}^{b}$ factors is dependent on the $L^{{\rm fIoU}}_{\rm NMS}$ and respectively.
@@ 3 Dataset description and network training
In this section, we present the main properties of the SDC1 data along with the expected products and the associated metrics. A complete description of the SDC1 challenge can be found in Bonaldi et al. (2021), while the underlying T-RECS simulation is detailed in Bonaldi et al. (2019). We also present the preprocessing of the data to construct our training sample. From this, we describe our best-performing network architecture and specify the corresponding setup and hyperparameters for our detector.
3.1 Subchallenge definition
@@
The SDC1 is a source detection and characterization task in simulated SKA-like data products (Sect. 1) that comprises nine 4GB images (three frequencies, with three integration times each) of the same field. The SDC1 is only modestly challenging regarding data volume, especially compared to the SDC2 with its 1TB data cube. Still, it represents significant challenges for detection methods in many other aspects. All the images have an identical pixel size of 32768x32768. As the frequency increases, the angular resolution improves while the field of view reduces. Therefore, images at different frequencies only partially overlap, meaning that the problem to solve varies with the position in the field. In addition, the number of detectable sources varies significantly with the integration time and frequency (Table 2 in Bonaldi et al., 2021). All the images are considered noise-limited, even at the highest 1000h integration time. As a benchmark for our YOLO-CIANNA method, we used only the 560 MHz - 1000h image due to its wider field of view, higher total number of sources, and higher source density per square pixel. This image is a fair example of a typical astronomical detection context with a high dynamic range and a high density of small sources with occasional blending. While our method could technically work for the other SDC1 images, it would not be more informative regarding its capabilities. The consequences and limits of this choice are further discussed in Sect. 5.2.1.
@@
In the original challenge, participating teams were provided a True catalog for a small fraction of the image. This was supposed to facilitate method development and tuning but also allow training of supervised ML approaches. At the time, the full underlying True catalog was unavailable, and the teams had to submit their result to a remote scoring service. After the challenge, the organizers released the full True catalog, the scoring code in open source [^2] https://gitlab.com/ska-telescope/sdc/ska-sdc (Clarke & Collinson, 2021), and the submitted source catalogs from the participating teams. In this study, we use only the training catalog to constrain our detector so our results can be compared to those published in the context of the challenge. We then use the full True catalog only to present an in-depth analysis of our trained detector performances. Since the scorer allows individual image scoring (for all frequencies and exposure), we can produce a detection catalog for the 560 MHz - 1000h image that is directly comparable to other team submissions during and after the challenge.
@@ 3.2 Image and source catalog description
The 560 MHz - 1000h image has a field of view of $5.5$ square degrees, which contains the primary beam out to the first null. It has a full size of 32768 square pixels, a pixel size of $0.6$ arcsec, and an imaging resolution in the Gaussian approximation of $1.5$ arcsec full width at half-maximum. A subfield of 512 square pixels from this image is presented in Fig. 1. The simulated image contains no systematic instrumental effects like calibration, pointing, or deconvolution errors, making it unrealistically clean (Sect. 5.3). The image is not primary-beam corrected, but the corresponding primary beam image is available as ancillary data. With this setup, the instrument sensitivity decreases on the edges, but it preserves a uniform noise level over the full field. The thermal noise level reported for this image in Bonaldi et al. (2021) is $2.55{\times}10^{-7}$ $\mbox{Jy}\,\mbox{beam}^{-1}$; however, a direct estimate of the noise using the provided image is about $3.45{\times}10^{-7}$ $\mbox{Jy}\,\mbox{beam}^{-1}$. We use this second value for the rest of the paper when necessary. While we could convert the image to a beam-corrected one, it is better for the detector to work on the constant noise image and to detect sources from their apparent brightness (Sect. 3.5). The apparent flux of a source can be obtained by multiplying the flux from the True catalog by the interpolated primary beam value at the central position of the source.
[…]
@@ 3.3 SDC1 scoring metric
The first element of the scorer is a match criterion. Due to source density, relying only on the central positions of the sources for matching predictions with the underlying True catalog is likely to result in many false positives. The SDC1 scorer uses a combination of the position, the size, and the predicted flux accuracies to represent a global matching score defined as
$E_{\rm tot}$ $=\sqrt{E^{2}_{\textrm{pos}}+E^{2}_{\textrm{size}}+E^{2}_{\textrm{flux}}}\text{, with}$ (15)
$E_{\textrm{pos}}$ $\propto\sqrt{(x-\hat{x})^{2}+(y-\hat{y})^{2}}/\hat{S}^{\prime}\text{,}$ (16)
$E_{\textrm{size}}$ $\propto\left|S-\hat{S}\right|/\hat{S}^{\prime}\text{, and}$ (17)
$E_{\textrm{flux}}$ $\propto\left|f-\hat{f}\right|/\hat{f},$ (18)
where the true values from the catalog are indicated with a hat, ($x,y$) is the central coordinates in pixels corresponding to (RA, Dec), $S$ is the average value of Bmaj and Bmin, $S^{\prime}$ is the largest axis convolved with the synthesized-beam, and $f$ is the source intrinsic flux. We note that the difference between $S$ and $S^{\prime}$ is present in the scorer code but not specified in either Bonaldi et al. (2021) or in the challenge documentation. To prevent false detections due to the high source density, the scorer uses a strict match threshold value for $E_{\rm tot}$. Each subpart of the error is normalized to a value considered representative of a $1\sigma$ error. In the scorer, the normalization coefficients are set to $0.93$ for the position, $0.36$ for the flux, and $4.38$ for the size, all using the catalog units defining a global $1\sigma$ error. These individual limits were obtained by fitting the distribution of the corresponding values on the combination of all the submitted catalogs at the challenge time. A match is defined if $E_{\rm tot}<5\sigma$, which was optimized to reduce the average random association chance of all submitted catalogs below 10%. We note that the scorer distinguishes false detection into two categories: either “False” if there is no target source closer than $1.5\times S^{\prime}$ or “bad” if it passes this distance limit but has a too-high $E_{\rm tot}$.
The list of identified matches is then converted to a global score that captures the detection and characterization quality. The SDC1 scorer attributes a score of up to one point for each match depending on its characterization quality and penalizes every false detection with a strict minus one point. The characterization evaluation is decomposed into seven individual subscores that all respect the following scoring rule (based on the scorer code) for a single source
@@
where $j$ represents the subscore part and $i$ a single source from the match list, $T^{j}$ is a threshold for this subscore part, and finally, $E_{i}^{j}$ and $s_{i}^{j}$ correspond to the error term and the final subscore part for this source. We list all the subscore parts and their corresponding $E^{j}$ function and $T^{j}$ values in Table 1, and we illustrate the typical score as a function of the error regarding a given threshold value in Fig. 6.
We note that all subpart error functions are based on a relative error, which has an asymmetric behavior. When overestimating the value, the error can rise infinitely, while underestimating a strictly positive prediction will never lower the relative error below -1. The issue is that these errors are associated with symmetric score response functions. Consequently, the score will be higher when underestimating the predictions, which is unlikely to naturally happen on quantities with intrinsic minimum values linked to instrumental limits like the flux, Bmaj, and Bmin. While the effect is minor for the subpart scoring, we note that for the matching criteria, it will result in excluding properly detected faint sources for which the noise caps the minimum predictable flux (Sects. 4.5 and 5.1).
The final score for a given source and the average subpart score for all sources are
$s_{i}=\frac{1}{7}\sum_{j}^{7}s_{i}^{j}\quad\text{ and }\quad\bar{s}^{j}=\sum_{i}^{N_{\textrm{match}}}s_{i}^{j}.$ (20)
@@
$\bar{s}=\frac{1}{N_{\textrm{match}}}\sum_{i}^{N_{\textrm{match}}}{s_{i}},$ (21)
@@
Table 1: Error functions and threshold values for all subscores.
Subpart Error function $E^{j}$ Threshold $T^{j}$
@@
Major axis |BmajBmaj^|/$\left|\textrm{Bmaj}-\hat{\textrm{Bmaj}}\right|/\hat{\textrm{Bmaj}}$ 0.3
Minor axis |BminBmin^|/$\left|\textrm{Bmin}-\hat{\textrm{Bmin}}\right|/\hat{\textrm{Bmin}}$ 0.3
Position angle $\left|\textrm{PA}-\hat{\textrm{PA}}\right|$ 10.0